OpenStudy (anonymous):

Find the difference. 5wx2 - 2wx2

OpenStudy (anonymous):

3wx^2

OpenStudy (anonymous):

ty... how do you do it? do you subtract the first but not the last?

OpenStudy (anonymous):

use FOIL

OpenStudy (anonymous):

?

OpenStudy (anonymous):

whats FOIL?

OpenStudy (anonymous):

first outside inside last

OpenStudy (anonymous):

there is no such thing as "foil" it is something dumb math teachers say

OpenStudy (anonymous):

well it works

OpenStudy (anonymous):

\[5wx^2 - 2wx^2\] is exactly the same as \(5-2\)

OpenStudy (anonymous):

satellite73 can u help me after plz

OpenStudy (anonymous):

since \[5-2=3\] you know \[5wx^2 - 2wx^2=3wx^2\]

OpenStudy (anonymous):

i said the answer already but thanx for explainng it

OpenStudy (anonymous):

there is no foil here tin-foll aluminium foil curses foiled again no such mathematical operation as foil

OpenStudy (anonymous):

lol hahaha omg but yess it is i learned it and it works with problems like these

OpenStudy (anonymous):

no actually this is called "combining like terms"

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

"foil" is a nmeonic devise for remembering that which has nothing whatsoever to do with this \[(a+b)(c+d)=ac+ad+bc+bd\]

OpenStudy (anonymous):

PENDAS is one way you can solve a problem... but only for multi step equations or 2 step equations

OpenStudy (anonymous):

yea you can use pemdas

OpenStudy (anonymous):

pemdas is another math teacher lie ignore it

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

there is no such thing as pembas,anymore than there is a foil

OpenStudy (anonymous):

it works so kids can learn it better

OpenStudy (anonymous):

no it only confuses them because it is a lie

OpenStudy (anonymous):

no it dones't confuse anybody it makes the problem easier

OpenStudy (jdoe0001):

\(\bf 5{\color{brown}{ wx^2}} - 2{\color{brown}{ wx^2}}\implies \square?{\color{brown}{ wx^2}} \)

OpenStudy (anonymous):

\[20\div 5\times 2\]

OpenStudy (anonymous):

its not... its pretty good i got 101% on things with PEMDAS... all trough middle school (5th, 6th, and 7th grade)

OpenStudy (anonymous):

according to "pemdas" i do the multiplication first and get \[20\div 10=2\] but that is way wrong ignore pemdas

OpenStudy (anonymous):

another bad mnemonic device to confuse students

OpenStudy (anonymous):

no i used it threw out middle school so it does work because i made 100%

OpenStudy (anonymous):

no you divide first... m and d share... so in that case 20 /5 is 4 and 4 times 2 is 8... checked answer... correct. :)

geerky42 (geerky42):

math-geek is correct pe(md)(as) it doesnt say you have to do multiplication first, just either multiplication or division first, depending which's first by reading left to right, before add or subtract.

geerky42 (geerky42):

pemdas isn't lie, it's just that most students misunderstood it and they though they have to do multiplication first... @satellite73

OpenStudy (anonymous):

(13x2 + 5x) + 8x2=__

OpenStudy (anonymous):

anyone know?